OpenStudy (anonymous):

translate and solve. the product of seven and the difference between a number times negative two and two is negative forty two

OpenStudy (anonymous):

We have product of 7, and "difference between a number times two, and two" equals -42. Rewrite this as an equation, and let a be the unknown number: \[7 \times [(a \times -2) - 2] = -42\] divide by 7 on both sides: \[(a \times -2) - 2 = -6]\ add 2 two both sides: \[a \times -2 = -4]\ finally divide by -2 on both times to isolate a \[a = 2]\

OpenStudy (anonymous):

We have product of 7, and "difference between a number times two, and two" equals -42. Rewrite this as an equation, and let a be the unknown number: \[7 \times [(a \times -2) - 2] = -42\] divide by 7 on both sides: \[(a \times -2) - 2 = -42\] add 2 to both sides:

OpenStudy (anonymous):

(continued) \[a \times -2 = -4\] finally, divide by -2 on both sides to get a: \[a = 2\]

OpenStudy (anonymous):

what is the integral of x sin x dx